OpenStudy (anonymous):

suppose that $2300 in initially invested in an account at a fixed interest rate, compounded continuously. After 3 years, the amount of money in the account is $2672. Find the interest rate per year. The answer should be written as a percentage. please and thank you!!

ganeshie8 (ganeshie8):

\[M = P e ^{r.t}\]

ganeshie8 (ganeshie8):

\[2672 = 2300 e ^{r . 3}\]

OpenStudy (anonymous):

oh okay, then what do i do??

ganeshie8 (ganeshie8):

divide both sides by 2300

OpenStudy (anonymous):

okay then?

ganeshie8 (ganeshie8):

\[2672/2300 = e ^{3r}\]

ganeshie8 (ganeshie8):

take natural log both sides

ganeshie8 (ganeshie8):

\[\ln (2672/2300) =3r\]

ganeshie8 (ganeshie8):

r = 5 %

OpenStudy (anonymous):

thats per year??

ganeshie8 (ganeshie8):

correct.. r = 3 years right... so its 5 % per year

OpenStudy (anonymous):

oh okay thank you so muchhh!

ganeshie8 (ganeshie8):

spend some time on "e" in your free time.. its an amazing number